OpenStudy (anonymous):

a solution contains 0.15 liters of sulfric acid. If this represents 10% of the total amount, find the total amount. Give the answer to one decimal place, if necessary

OpenStudy (anonymous):

0.15*100/10=1.5litres

OpenStudy (anonymous):

total is 1.5 liters?? Why so small?

OpenStudy (amistre64):

.15 is 10%; so divide it by 10 to see it at 1% .15/10 = 015; now to see it at 100% we just multiply it by 100 .015 * 100 = 1.5

OpenStudy (anonymous):

Ohhh! Now I see :) Thankyou :)